OpenStudy (anonymous):

If a=12 and b= 12 solve for c by using the Pythagorean Theorem. Next, give the exact answer in simplified radical form. Finally, give the answer rounded to the nearest hundredth.

OpenStudy (anonymous):

pythagorean theorm = a^2 + b^2 = c^2 therefore 12^2 + 12^2 = 288 to find c we square root that answer to get 16.970 or 16.97 to the nearest hundreth in simplified radical form that is simply 12sqrt2 Hope this helps! Cheers! ^.^ Sophia